determine if set of vectors are linearly independent v1={4,1,2}, v2={-3,0,1} and v3={1,2,1}


1
Expert's answer
2021-06-07T19:37:23-0400

Let us determine the determinant of the matrix consisting of coordinates of vectors:

"d = \begin{vmatrix}
 4 & 1 & 2 \\
 -3 & 0 & 1 \\
1 & 2 & 1
\end{vmatrix} = 4\begin{vmatrix}
 0 & 1 \\
 2 & 1
\end{vmatrix} - 1 \begin{vmatrix}
 -3 & 1 \\
 1 & 1
\end{vmatrix} + 2 \begin{vmatrix}
 -3 & 0 \\
 1 & 2
\end{vmatrix}" ,

"d = 4\cdot(0\cdot1 -2\cdot1 ) -1\cdot(-3\cdot1 -1\cdot 1)+2\cdot(-3\cdot2 - 1\cdot0), \\
d = -16."

The determinant is not 0, so the system of vectors is linearly independent.